OpenStudy (671grains):

help got the aster to the first part of the question need help with the second part

OpenStudy (671grains):

second part is Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal the answer to the first part isA(t)=39145(1+0.03/12)^12t

OpenStudy (671grains):

someone please help if you can this is really late

OpenStudy (anonymous):

set A(t) = 10,000 and solve for t

OpenStudy (anonymous):

\[\log _{(1+\frac{ 0.03 }{ 12 })}\frac{ 10000 }{ 39145 }\frac{ 1 }{ 12 }=t\]

OpenStudy (671grains):

ok one sec let me do the math really quick

OpenStudy (anonymous):

for the logarithm you can use change of base formula \[\frac{ \ln(\frac{ 10000 }{ 39145 }\frac{ 1 }{ 12 }) }{ \ln(1+\frac{ .03 }{ 12 }) }=t\]

OpenStudy (anonymous):

the equation should also be negative I am not getting a reasonable number right now though

OpenStudy (671grains):

so is it 0.239? i don't think I did it right but I am not sure

OpenStudy (anonymous):

my answer comes out as 1542 years

OpenStudy (anonymous):

or months

OpenStudy (anonymous):

so 128 years

OpenStudy (anonymous):

I don't think this is correct

OpenStudy (671grains):

so the 0.239 would be -0.239 if it is correct

OpenStudy (anonymous):

I'm not getting the right answer either I don't think. My equation should also have a negative vvalue for t

OpenStudy (671grains):

ya neither do I

OpenStudy (anonymous):

Not sure about this once. I don't see where else there could be an error The equation should be\[\frac{ \ln(\frac{ (12)39145 }{ 10000 }) }{ \ln(1+\frac{ .03 }{ 12 }) }=t\]

OpenStudy (671grains):

n is 12 so if it was 1n wouldn't be 1(12) which is 12

OpenStudy (anonymous):

Im not sure what n is. Maybe paste the whole problem

OpenStudy (671grains):

It won't let me